OpenStudy (anonymous):

∫((x^2 -x-5)/(x-3)) dx Please show work

sam (.sam.):

∫[(x^(2)-x-5)/(x-3),x] Partial fraction x+2+(1)/(x-3) let u=x-3 dx=du ln(|u|)+C (x^(2))/(2)+2x+ln(|(x-3)|)+C

OpenStudy (anonymous):

∫[(x^(2)-x-5)/(x-3),x] whered you get the x at the end here?

sam (.sam.):

thats dx, wrote it wrongly

OpenStudy (anonymous):

oh okay, thank you for your help!

sam (.sam.):

welcome

OpenStudy (anonymous):

sam, in your solution, the final part, what is the x^2 over? all of it? or only 2?

OpenStudy (anonymous):

sam?